Ellix Company manufactures two models of ultra-high fidelity speakers—the X200 model and the X99 model. Data regarding the two products follow:

Product Direct Labor-Hours Annual
Production
Total Direct
Labor-Hours
X200 0.4 DLHs per unit 25,000 units 10,000 DLHs
X99 0.6 DLHs per unit 45,000 units 27,000 DLHs
37,000 DLHs

Additional information about the company follows:

a. Model X200 requires $33 in direct materials per unit, and model X99 requires $23.

b. The direct labor workers are paid $40 per hour.

c. The company has always used direct labor-hours as the base for applying manufacturing overhead cost to products.

d. Model X200 is more complex to manufacture than model X99 and requires the use of special equipment.

e. Because of the special work required in (d) above, the company is considering the use of activity-based absorption costing to apply manufacturing overhead cost to products. Three activity cost pools have been identified as follows:

Estimated Total Activity
Activity Cost Pool Activity Measure Estimated
Total Cost
X200 X99 Total
Machine setups Number of setups $ 205,200 108 162 270
Special processing Machine-hours 313,100 1,010 0 1,010
General factory Direct labor-hours 2,220,000 10,000 27,000 37,000
$ 2,738,300

Required:

1. Assume that the company continues to use direct labor-hours as the base for applying overhead cost to products.

a. Compute the plantwide predetermined overhead rate.

b. Compute the unit product cost of each model.

c. Assume that the company decides to use activity-based absorption costing to apply overhead cost to products.

d. Compute the activity rate for each activity cost pool and determine the amount of overhead cost that would be applied to each model using the activity-based approach.

e. Compute the unit product cost of each model.
Answer #1
Plantwide rate = Estimated Overheads/Estimated Labor hours
=2738300/37000
74.00810811 per hour
Calculation of Unit product Cost
X200 X99
Direct material per unit 33 23
Direct labor per unit 16 24
Overhead cost per unit 29.60324324 44.40486486
Unit product cost 78.60324324 91.40486486
Estimated Total Activity Applied = Rate*Driver consumed
Activity Cost Pool Activity Measure Estimated X200 X99 Total Activity rate = Estimated Overheads/Total Activity X200 X99
Total Cost
Machine setups Number of setups 205,200 108 162 270 760 82080 123120
Special processing Machine-hours 313,100 1,010 0 1,010 310 313100 0
General factory Direct labor-hours 2,220,000 10,000 27,000 37,000 60 600000 1620000
Total 2,738,300 995180 1743120
Calculation of Unit product Cost
X200 X99
Direct material per unit 33 23
Direct labor per unit 16 24
Overhead cost per unit 39.8072 38.736
Unit product cost 88.8072 85.736

А 1 Plantwide rate = Estimated Overheads/Estimated Labor hours 2 =2738300/37000 3 =2738300/37000 per hour 4 5 Calculation of